Consider a die with 2 red sides, 2 green sides, and 2 blue sides. Roll the die 5 times, and let X denote the number of times that the die has a red result.

Flip a coin 5 times, and let Y denote the number of times that the coin shows “heads."

a. Find E(X?Y).

b. Find Var?(X?Y).
Answer #1

Here X can take values 0, 1, 2, 3, 4 and 5. And Y can take values 0 ,1 ,2, 3, 4, 5. Since roll is independent from other so X has binomial distribution with parameter n=5 and p=2/ 6 = 1/3. The pdf of X is

P(X=x)=\binom{5}{x}\left ( \frac{1}{3} \right )^{x}\left ( \frac{2}{3} \right )^{5-x},x=0,1,2,3,4,5

Following table shows the probabilites of X, rounding to 4 decimal places:

X P(X=x)
0 0.1317
1 0.3292
2 0.3292
3 0.1646
4 0.0412
5 0.0041

-----------------

Here X has binomial distribution with parameter n=5 and p=0.5. The pdf of Y is

P(Y=y)=\binom{5}{y}\left ( 0.5 \right )^{y}\left ( 0.5 \right )^{5-y},y=0,1,2,3,4,5

Following table shows the probabilites of Y, rounding to 4 decimal places:

Y P(Y=y)
0 0.0313
1 0.1563
2 0.3125
3 0.3125
4 0.1563
5 0.0313

Let W=X-Y. Since X and Y are independent so  

P(W=w)= P(X=x)P(Y=y)

Following table shows the calculations for W:

X P(X=x) Y P(Y=y) W=X-Y P(W=w)
0 0.1317 0 0.0313 0 0.00412221
0 0.1317 1 0.1563 -1 0.02058471
0 0.1317 2 0.3125 -2 0.04115625
0 0.1317 3 0.3125 -3 0.04115625
0 0.1317 4 0.1563 -4 0.02058471
0 0.1317 5 0.0313 -5 0.00412221
1 0.3292 0 0.0313 1 0.01030396
1 0.3292 1 0.1563 0 0.05145396
1 0.3292 2 0.3125 -1 0.102875
1 0.3292 3 0.3125 -2 0.102875
1 0.3292 4 0.1563 -3 0.05145396
1 0.3292 5 0.0313 -4 0.01030396
2 0.3292 0 0.0313 2 0.01030396
2 0.3292 1 0.1563 1 0.05145396
2 0.3292 2 0.3125 0 0.102875
2 0.3292 3 0.3125 -1 0.102875
2 0.3292 4 0.1563 -2 0.05145396
2 0.3292 5 0.0313 -3 0.01030396
3 0.1646 0 0.0313 3 0.00515198
3 0.1646 1 0.1563 2 0.02572698
3 0.1646 2 0.3125 1 0.0514375
3 0.1646 3 0.3125 0 0.0514375
3 0.1646 4 0.1563 -1 0.02572698
3 0.1646 5 0.0313 -2 0.00515198
4 0.0412 0 0.0313 4 0.00128956
4 0.0412 1 0.1563 3 0.00643956
4 0.0412 2 0.3125 2 0.012875
4 0.0412 3 0.3125 1 0.012875
4 0.0412 4 0.1563 0 0.00643956
4 0.0412 5 0.0313 -1 0.00128956
5 0.0041 0 0.0313 5 0.00012833
5 0.0041 1 0.1563 4 0.00064083
5 0.0041 2 0.3125 3 0.00128125
5 0.0041 3 0.3125 2 0.00128125
5 0.0041 4 0.1563 1 0.00064083
5 0.0041 5 0.0313 0 0.00012833

Following table shows the pdf of W:

W P(W=w)
-5 0.00412221
-4 0.03088867
-3 0.10291417
-2 0.20063719
-1 0.25335125
0 0.21645656
1 0.12671125
2 0.05018719
3 0.01287279
4 0.00193039
5 0.00012833

(a-b)

Following table shows the calculations for mean and variance of W:

W P(W=w) wP(W=w) w^2P(W=w)
-5 0.00412221 -0.02061105 0.10305525
-4 0.03088867 -0.12355468 0.49421872
-3 0.10291417 -0.30874251 0.92622753
-2 0.20063719 -0.40127438 0.80254876
-1 0.25335125 -0.25335125 0.25335125
0 0.21645656 0 0
1 0.12671125 0.12671125 0.12671125
2 0.05018719 0.10037438 0.20074876
3 0.01287279 0.03861837 0.11585511
4 0.00193039 0.00772156 0.03088624
5 0.00012833 0.00064165 0.00320825
Total -0.83346666 3.05681112

So

E(X - Y)E(W)-w P(W-)-0.83346666

Var(X-Y)=Var(W)=\sum w^{2}P(W=w)-\left [ \sum wP(W=w) \right ]^{2}=2.362144447